OpenStudy (anonymous):

x^4/x^16 i dont know 12^10/12^2 12^8 5n^14/20n^7 i dont know 6^8/6^8 1 a^9b^5/a^3b^15 i dont know /=divided by ASAP HELP PLZ

ganeshie8 (ganeshie8):

use quotient of power property

ganeshie8 (ganeshie8):

\(\large \frac{a^m}{a^n} = a^{m-n}\)

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

m minus n?

ganeshie8 (ganeshie8):

\(\huge \frac{a^m}{a^n} = a^{m-n}\)

ganeshie8 (ganeshie8):

yup! u simply subtract the exponents

ganeshie8 (ganeshie8):

\(\huge \frac{x^4}{x^{16}} = ?\)

OpenStudy (anonymous):

wouldnt that be negative?

ganeshie8 (ganeshie8):

yes :)

OpenStudy (anonymous):

but i dont have negative as a choice

ganeshie8 (ganeshie8):

\(\huge \frac{x^4}{x^16} = x^{4-16}\)

ganeshie8 (ganeshie8):

\(\huge \frac{x^4}{x^16} = x^{4-16} = x^{-12}\)

ganeshie8 (ganeshie8):

wat choices do u have

OpenStudy (anonymous):

x12 1/x^12

ganeshie8 (ganeshie8):

\(\huge \frac{x^4}{x^16} = x^{4-16} = x^{-12} = \frac{1}{x^{12}}\)

ganeshie8 (ganeshie8):

answer for ur next question is correct

OpenStudy (anonymous):

a^9b^5/a^3b^15?

ganeshie8 (ganeshie8):

\(\huge \frac{a^9b^5}{a^3b^{15}}\)

ganeshie8 (ganeshie8):

\(\huge a^{9-3} ~ b^{5-15}\)

ganeshie8 (ganeshie8):

\(\huge a^{6} ~ b^{-10}\)

ganeshie8 (ganeshie8):

\(\huge\frac{a^{6}}{ b^{10}}\)

OpenStudy (anonymous):

5n^14/20n^7 would it be the second one? n^2/4 n^7/4

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

Thanks!!!!!

ganeshie8 (ganeshie8):

np:)

OpenStudy (anonymous):

another question... im a bit confused on it sorry 8b^3/20b^3